# Incorporating Physical Priors into Weakly-Supervised Anomaly Detection

### v1

@@ V Conclusions
We have introduced a Prior-Assisted Weak Supervision (PAWS) approach to anomaly detection. Like previous weakly-supervised methods, PAWS is based on training a classifier to distinguish data from a background-only reference dataset. We impose structure on the function class of classifiers through a pre-training step based on a set of signal models. This trades off breadth for depth of sensitivity - we show that PAWS is able to improve the sensitivity of previous methods by a factor of 10–100 in signal cross section. While we require pre-specifying the signal class, it can (and in our case, does) contain a large number of models that would be infeasible to probe one at a time. One could scan, in steps consistent with the experimental resolution, over the parameters of the models. In addition to the computational challenge, this would also incur a large trial factor that PAWS does not pay. Furthermore, the sensitivity of PAWS is at the limit of what is possible for any method and so there is little to be lost compared to a dedicated approach.
There are a number of extensions of PAWS that would be exciting to explore in future work. One direction would be to push the dimensionality of the parameter space. In our numerical example, the parameter space was three-dimensional ($m_{X},m_{Y}$, and $\alpha$) in addition to the signal fraction $\mu$. Adding parameters would simply require adding features to the classifiers, but it may be advisable to not scan over them uniformly when building the parameterized classifier. Another direction would be to push the dimensionality of the feature space. In our numerical example, the feature space was six-dimensional. Preliminary studies suggest that the setup works well when using the full phase space Buhmann:2023acn , but the performance is so good that more data is required for training and testing. It would also be interesting to explore the sensitivity to signals that are similar to the ones used in the pre-training, but not exactly the same. We found empirically that the model does not extrapolate well to $m_{X}$ and $m_{Y}$ outside of the training dataset, but this is not conclusive because the jet substructure also changes significantly (e.g. the decay products are no longer contained within a single jet). While our example focused on resonant anomalies, we expect that the PAWS strategy would work in any scenario where data is compared to a background-only reference. Exploring various combinations of PAWS with other weakly-supervised approaches for signal sensitivity and background estimation could lead to further improvements. Ultimately, we expect that a variety of strategies with complementary assumptions and strengths will be required to achieve broad discovery potential at the LHC and beyond.

### v2

@@ V Conclusions
We have introduced a Prior-Assisted Weak Supervision (PAWS) approach to anomaly detection. Like previous weakly-supervised methods, PAWS is based on training a classifier to distinguish data from a background-only reference dataset. We impose structure on the function class of classifiers through a pre-training step based on a set of signal models. This trades off breadth for depth of sensitivity - we show that PAWS is able to improve the sensitivity of previous methods by a factor of 10–100 in signal cross section. While we require pre-specifying the signal class, it can (and in our case, does) contain a large number of models that would be infeasible to probe one at a time. One could scan, in steps consistent with the experimental resolution, over the parameters of the models. In addition to the computational challenge, this would also incur a large trial factor that PAWS does not pay since it uses pretraining and gradient descent. Furthermore, the sensitivity of PAWS is at the limit of what is possible for any method using the same input features and so there is little to be lost compared to a dedicated approach.
There are a number of extensions of PAWS that would be exciting to explore in future work. One direction would be to push the dimensionality of the parameter space. In our numerical example, the parameter space was three-dimensional ($m_{X},m_{Y}$, and $\alpha$) in addition to the signal fraction $\mu$. Adding parameters would simply require adding features to the classifiers, but it may be advisable to not scan over them uniformly when building the parameterized classifier. Another direction would be to push the dimensionality of the feature space. In our numerical example, the feature space was six-dimensional. Preliminary studies suggest that the setup works well when using the full phase space Buhmann:2023acn , but the performance is so good that more data is required for training and testing. It would also be interesting to explore the sensitivity to signals that are similar to the ones used in the pre-training, but not exactly the same. We found empirically that the model does not extrapolate well to $m_{X}$ and $m_{Y}$ outside of the training dataset, but this is not conclusive because the jet substructure also changes significantly (e.g. the decay products are no longer contained within a single jet). While our example focused on resonant anomalies, we expect that the PAWS strategy would work in any scenario where data is compared to a background-only reference. Combining PAWS with other weakly-supervised approaches for signal sensitivity and background estimation could lead to further improvements. Ultimately, we expect that a variety of strategies with complementary assumptions and strengths will be required to achieve broad discovery potential at the LHC and beyond.